OpenStudy (anonymous):

(8x^2+17x)+6x^2= explain/check answer

OpenStudy (danjs):

if you have terms to the same power like ax2 you can add them together

OpenStudy (danjs):

Ax^2 + Bx^2 = (A+B)x^2

OpenStudy (anonymous):

hold on let me get a calculator

OpenStudy (danjs):

eight plus 6

OpenStudy (anonymous):

what about 17???

OpenStudy (danjs):

\[8x^2 + 17x + 6x^2 = 14x^2 + 17x\]

OpenStudy (danjs):

you can combine terms with the same power

OpenStudy (anonymous):

would it be 31^5???

OpenStudy (danjs):

what does it = to 0?

OpenStudy (danjs):

or you just simplifying

OpenStudy (anonymous):

it just says to find the sum

OpenStudy (danjs):

then you are just simplifying

OpenStudy (jhannybean):

Yep.

OpenStudy (anonymous):

okay so was my answer right?

OpenStudy (danjs):

Just combine the terms that have the same powers on the x's

OpenStudy (anonymous):

so is it 31^5????

OpenStudy (danjs):

where did you get 31?

OpenStudy (anonymous):

i added all of the x's together

OpenStudy (danjs):

\[8x^2 + 6x^2 = 14x^2\]

OpenStudy (danjs):

you add the coefficients of the like terms, x^2

OpenStudy (danjs):

the powers do not add

OpenStudy (anonymous):

14x^2+17x right?

OpenStudy (danjs):

correct

OpenStudy (anonymous):

and so they multiply?

OpenStudy (danjs):

that is all you can do

OpenStudy (anonymous):

no i meant the exponents

OpenStudy (danjs):

the exponents multiply, if you have something like this

OpenStudy (anonymous):

ohh so if they are different i add?

OpenStudy (danjs):

\[[x^2]^2 = x ^{2*2} = x^4\]

OpenStudy (danjs):

A power to a power , you multiply the powers

OpenStudy (danjs):

You add the powers in this case....

OpenStudy (danjs):

\[x^2 * x^2 = x ^{2+2} = x^4\]

OpenStudy (anonymous):

oh okay i got it, thank you

OpenStudy (danjs):

we are just adding terms here, not multiplying them

OpenStudy (danjs):

so you just add the coefficients in front of the x^2 and keep the same power

OpenStudy (danjs):

\[8x^2 + 17x +6x^2 = [8x^2 + 6x^2] +7x = (8+6)x^2 +7x = 14x^2 + 7x\]

OpenStudy (danjs):

that is the summary